what is the integration of the following statement
(sqrt(1-4x^2) / x) dx

Grade:12th pass

1 Answers

Aditya Gupta
4 years ago
(sqrt(1-4x^2) / x)=(x*sqrt(1-4x^2) / x^2) now put y=1 – 4x^2 so that dy= – 8xdx
so we have int 4sqrty/(1-y) dy/-8 = – ½ int sqrt y/(1-y) dy
now put sqrt y= z or y = z^2 and the ques becomes too easy to solve